The printed matter on a 14 by 16 centimeter page of a book must cover 80 square centimeters. If all margins are to be the same width, how wide should they be?
Answer by checkley77(12844)

(14-2x)(16-2x)=80
224-32x-28x+4x62=80
4x^2-60x+224-80=0
4x^2-60x+144=0
4(x^2-15x+36)=0
4(x-12)(x-3)=0
x-3=0
x=3 ans. for the width of the margins.
Proof:
(14-2*3)(16-2*3)=80
(14-6)(16-6)=80
8*10=80
80=80